A bag contains 150 balls numbered 1 through 150. Three balls are drawn and placed on the table. The probability that the balls were drawn in increasing order can be expressed as a/b where a and b are co-prime positive integers. What is the value of a+b?

Grade:11

1 Answers

Shivam Dimri
10 years ago

well this is not so typical,

selection of any three balls will give 3 distinct numbers

hence suppose they are a,b,c

they are all distinct here remember, hence such selection will always allow us to arrange them in increasing order without thinking of what balls we selected and thats the best part!!!

probabilty = 1/3!

 answer = 1/6

so a + b = 7

!!

do reply..